Quantum Technology

Quantum Technology had $669,000 of retained earnings on December 31, 20X2. The company

paid common dividends of $35,500 in 20X2 and had retained earnings of $576,000 on December

31, 20X1. How much did Quantum Technology earn during 20X2, and what would earnings per

share be if 47,400 shares of common stock were outstanding?
Answer #1

Retained earnings, December 31, 20X2 ........................ $669,000 

Less: Retained earnings, December 31, 20X1............... (576,000) 

Change in retained earnings.................................. $93,000 

Add: Common stock dividends...................................... 35,500 

Earnings available to common stockholders................. $128,500 


*Earnings per share = $128,500/47,400 = $2.71 per share 

answered by: Mazida Rahman
Answer #2

SOLUTION : 


Earning in 20X2

= RE in 20X2 - RE in 20X1 + dividends paid in 20X2

= 669000 - 576000 + 35500

= 128500 ($) (ANSWER)


Earning per share 

= Earnings / Number of shares

= 128500 / 47400

= 2.71 ($) (ANSWER).

answered by: Tulsiram Garg